OpenStudy (anonymous):

I need help explaining the solution of this problem... Jessie just finished the probation period at work and got a $0.50 per hour raise and 4 more hours of work per week. Jessie's friend scoffed, "That's only $2 more per week! Who cares?" Why is Jessie's friend wrong about the change in Jessie's weekly pay? What more do we need to know about Jessie's work situation to calculate the actual change in the pay? How is this related to the product rule in calculus class?

OpenStudy (camper4834):

He is wrong cuz jessie is earning more than $2 per week. What we need to know is how many hours he was working before as that accounts for it. Product rule in calculus? Gotta look that up. Brb

OpenStudy (camper4834):

Oh product rule of derivatives?

OpenStudy (anonymous):

correct. I have one idea of the 'formula' but I can't figure it out how to relate to the product rule. x=original hours worked y=original new per hour xy=overall pay new pay= (x+4)(y+0.5) = xy+.5x+4y+2 increase= 0.5x+4y+2

OpenStudy (camper4834):

i guess most of your question was really that last part huh?

OpenStudy (anonymous):

Kind of I wanted to make sure my formula was correct.

OpenStudy (camper4834):

Your formula looks good, im not sure what they are looking for. sorry..